Jane wants to purchase a bond with a face value of 9,750 and a bond rate of 6% per year, payable quarterly. The bond has a remaining life of 4 years. If Jane wants to earn at least 6% per year compounded quarterly, what is the maximum price she would be willing to pay to purchase the bond?
Answer #1

i = 6% / 4 = 1.5% per quarter

t = 4 * 4 = 16 quarters

coupon payment = 6% / 4 * 9750 = 146.25

Present value of bond = 146.25 * (P/A,1.5%,16) + 9750 * (P/F,1.5%,16)

= 146.25 *14.131264 + 9750 *0.788031

= 9749.999 ~ 9750 (Nearest Dollar)

We could have given same value as par value of bond, as coupon rate of bond is equal to interest rate. I gave calculations just to show the steps